# QCD Phase Diagram and the Finite Volume Fireball: A Model Study

### v1

@@ Abstract
Using the Nambu Jona-Lasinio model, an effective theory for the chiral transition of quantum chromodynamics (QCD), we have studied the effect of the finite volume of the fireball on the transition line at finite temperature and density using the MIT boundary condition to mimic the condition that the system is deconfined inside. The shift of the transition temperature for finite volume and finite volume effect on number density and its susceptibilities are studied. The volume effects should be considered when looking for signatures of the phase diagram in experiments.
@@ I Introduction
The bulk property of the strongly interacting matter has been a subject of detailed experimental investigation in recent times [1, 2, 3, 4]. Strongly interacting matter is expected to manifest itself in a deconfined, chiral symmetry-restored phase, called the quark-gluon plasma (QGP), at very high temperatures [5, 6]. Experimentally, this phase has been explored in the study of the medium created in the collision of heavy ions at ultrarelativistic speeds, in particular in the Relativistic Heavy Ion Collider (RHIC) in the USA [3] and the Large Hadron Collider (LHC) in CERN, Geneva [1]. The phase structure at high number density is being explored in the Beam Energy Scan (BES) experiments in RHIC [4] and will be further explored in future experiments in NICA, JINR [7]. The knowledge about the phases at high density is complemented by the data from neutron stars and their mergers [8]. On the theory side, the properties of the deconfined state at high temperatures are well-understood from large-scale numerical studies of quantum chromodynamics (QCD), the theory of strongly interacting matter, using lattice discretization [5]. Lattice studies are more difficult at nonzero baryon number chemical potential ($\mu_{\scriptscriptstyle B}$); however, the equation of state and baryon number susceptibilities have been calculated at large temperatures and moderate values of $\mu_{\scriptscriptstyle B}$. At higher values of $\mu_{\scriptscriptstyle B}$, or lower temperatures, one has to rely on QCD-based models [9, 10].
@@
The theoretical studies, in particular the lattice-discretized QCD (LQCD) studies, are for static equilibrated infinite volume systems. On the other hand, the medium created in RHIC and LHC experiments is rapidly evolving, and the system size is not too large. In particular, in the BES studies the system dimension can be a few fermi, not too large compared to the typical correlation length in strong interaction physics. The current experimental efforts are to study QGP in smaller systems. All this makes an understanding of the volume effect important.
@@
In this paper, the aim is to get an idea of the effect of the finite volume on the phase diagram for a static system, with the effects of the dynamic nature left for a future study. We study the phase diagram using the Nambu Jona-Lasinio (NJL) model [9], which is an effective model expected to capture the physics related to the chiral symmetry, in particular, the spontaneous chiral symmetry breaking. Since the finite temperature crossover is associated with the melting of the chiral condensate, the NJL model is expected to capture the basic features of the crossover.
@@
In a study of the finite volume effects it is essential that realistic boundary conditions are put in for the system. The actual boundary of the medium created in the heavy ion collision experiments is of complicated geometry. We take the simplified MIT boundary condition [11] (see [12] for a pedagogic discussion). While simple, the MIT boundary condition captures the essential physics that inside the boundary we have a deconfined medium.
@@
The effect of the finite volume on the chiral transition has been studied before [13, 14, 15, 16, 17, 18, 19, 20]. These studies have used the anti-periodic (APBC) and periodic (PBC) boundary conditions, or have put in a lower cutoff in momentum. However, these boundary conditions do not mimic a finite-volume fireball; they are boundary conditions used in theoretical studies to mimic an infinite volume system within a finite-volume setup, e.g., in numerical lattice studies. In the Appendix B we compare the MIT boundary condition results with those obtained with other geometries. The MIT boundary condition has been used before, for studying the chiral crossover at $\mu_{\scriptscriptstyle B}=0$ on a sphere [21], and for a rotating fermionic system in a cylinder [22, 23]. Here we use it to study the phase diagram in the $\left(T,\,\mu_{\scriptscriptstyle B}\right)$ plane and the baryon number susceptibilities.
@@ II Formalism and the boundary condition
The NJL model is an effective model of the spontaneous chiral symmetry breaking of QCD [24, 9]. The degrees of freedom are only quarks; the gluonic interactions are included in four-fermion interaction terms, chosen such that the Lagrangian is chirally symmetric.
@@
The four-fermion term in $\mathcal{L}_{\scriptscriptstyle\rm NJL}$ makes the theory non-renormalizable. Hence, a suitable regularization method has to be specified, and the results will depend on the regularization parameter. We have used Schwinger’s proper time regularization [24, 25], with the cutoff $\tau_{\rm\scriptscriptstyle UV}=1/\Lambda_{\rm\scriptscriptstyle UV}^{2}$. Therefore, our model has three parameters: $m,G$ and the UV momentum cutoff $\Lambda_{\rm\scriptscriptstyle UV}$. These parameters can be chosen by specifying the pion mass, the pion decay constant and either the chiral condensate or the chiral transition temperature ($T_{c}$). Some suitable parameter sets, and how they affect the chiral crossover temperature, have been discussed in Ref. [25]. The phase diagram can be qualitatively different for the different parameter sets. In particular, for some choices of parameters one gets a critical point and a first-order transition at large quark number chemical potential $\mu_{q}=\dfrac{\mu_{\scriptscriptstyle B}}{3}$, while for some other choices, the transition is a crossover in the whole $\left(T,\,\mu_{q}\right)$ plane. We have chosen the following parameter set:
@@
This parameter set was already explored in Ref. [25]. It has a first-order transition line at large $\mu_{q}$, ending at a critical point at $\mu_{q,c}\approx$ 350 MeV. At $\mu_{q}=0$, it shows a chiral transition at $T_{\chi}\approx$ 186 MeV. Both the $T_{\chi}$ and the $\mu_{q,c}$ are larger than the expected values for QCD. However, here our aim is not to tune the absolute value of these parameters but to look at their variation with the system size.
We will be studying the system at finite temperature $T$ and quark number chemical potential $\mu_{q}$. $\mu_{q}$ is introduced by adding $\mu_{q}\psi^{\dagger}\psi$ to Eq. (1). To get the thermodynamic potential, the integral over $p_{0}$ in Eq. (4) has to be turned into a fermionic Matsubara sum. For finite volume, the integral over spatial momenta, $\dfrac{d^{3}p}{(2\pi)^{3}}$, will be replaced by a sum over the momentum modes allowed by the boundary condition. For a study of the effect of finite volume, imposing a proper boundary condition is important. As we discussed in Section I, the exact boundary condition for the medium created in RHIC or LHC is complicated; since it is time-dependent and also changes from event to event. Since we are only interested in getting an idea about how different the finite volume system can be from an infinite volume system, we will take a simple spherical geometry. The important physics that we want to capture is that the deconfined QGP is only contained within this geometry. For this, we impose the MIT boundary condition, which implies that the normal component of the quark current becomes zero at the boundary. More precisely, we impose the boundary condition [11, 12],
@@
Effects of finite volume on the phase diagram have been explored in the literature in the past. Sometimes the antiperiodic or the periodic boundary conditions have been used for this purpose [15, 16, 26, 19]. Note that these boundary conditions do not capture the finite volume fireball in any way: these are boundary conditions usually imposed in theoretical studies (e.g. numerical studies on the lattice) when the aim is to study the infinite volume system. Besides being easy to implement, the advantage of these boundary conditions is that they show very small deviations from the infinite volume system, and are therefore ideal for studies aiming for the thermodynamic limit. Our purpose here is to estimate the size of the deviation from the infinite volume results due to the finite volume fireball. We do not think (anti)periodic boundary conditions are suitable for this purpose. While the boundary condition Eq. (5), Eq. (6) also do not capture the whole complexity of the finite fireball, we believe it captures the most essential aspects of it and gives a more realistic estimation of the size of the finite volume effects. In Appendix B we compare our results with results with an antiperiodic geometry for a similar system size. We indeed find that the volume effects are much smaller in the antiperiodic geometries.
@@ III Chiral transition at finite volume
We explore the chiral transition line for Eq. (3), and its shift with volume, for the MIT boundary condition. At $\mu_{q}=0$ the chiral symmetry restoration is a crossover. In Fig. 1, we show the constituent mass, $M$ (Eq. (10)), which is just a constant shift of $-\langle\bar{\psi}\psi\rangle$ and has the same information as $\langle\bar{\psi}\psi\rangle$. We see a rapid but smooth decrease in $M$ around $T\approx$ 190 MeV, indicating a crossover. At $T=0$, decreasing the volume to $R=5$ fm decreases the constituent mass by $<5\%$ from the infinite volume value. At high temperatures $>200$ MeV the chiral symmetry is restored and the chiral condensate becomes small. Therefore $M$ becomes small, approaching $m$. In small volume systems, the crossover region becomes somewhat smoother, but no qualitative change from the infinite volume case is seen in systems as small as a 5 fm sphere. This can be further seen in the right panel of Fig. 1, where the derivative $dM/dT$ is shown. To get this plot, we used a cubic spline interpolation of the $M$ vs $T$ data of the left plot. The plot shows a systematic but moderate reduction in the peak height and broadening of the peak, as the volume is reduced. The peak position is an estimate of the crossover temperature $T_{c}$, which changes very slowly with the radius of the sphere, $R$ (a $7$ MeV shift from infinite volume to $R=5$ fm).
@@
Figure 1: (Left) Constituent quark mass $M$ as a function of temperature ($T$) at $\mu$ = 0. (Right) The derivative $dM/dT$. The legend shows the peak position of the $dM/dT$ curves, which can be taken as an estimate of $T_{c}$.
Instead of $dM/dT$, it is more common to look at the chiral susceptibility $\chi_{m}=-\frac{\textstyle\partial\langle\bar{\psi}\psi\rangle}{\textstyle\partial m}$. The chiral susceptibility shows a peak at the chiral symmetry restoration point. The peak position is close to the peak of $dM/dT$, though not exactly identical at the crossover region, and the volume dependence is similar to that seen in Fig. 1. Here we will use the peak of $dM/dT$ as our estimate for $T_{c}$. Of course, in the case of a true phase transition (as opposed to a crossover), all these observables will show the same transition temperature.
@@
Now we introduce a nonzero quark chemical potential $\mu_{q}$. For small values of $\mu_{q}$, the qualitative behavior is very similar to what is seen for $\mu=0$. In Fig. 2, we show $M$ as a function of $T$ at different volumes for three values of $\mu_{q}$. The temperature and volume dependence of $M$ at $\mu_{q}$=100 MeV is very similar to that at $\mu$=0, as a comparison of the left panel of Fig. 2 with Fig. 1 shows. As $\mu_{q}$ increases, the crossover shifts to smaller values of temperature, as expected. Also, we find a stiffening of the transition at larger volumes: the smoothing of the transition at small volumes is enhanced. At $\mu_{q}$=350 MeV, shown in the right panel of Fig. 2, we find a discontinuity, indicating a first-order transition, at infinite volume. The first-order nature of the transition can be further ascertained by looking at $\Omega$, Fig. 3. At the transition region, a two-minimum structure is seen, which is the hallmark of first-order transition, and indicates phase coexistence. At this $\mu_{q}$ the transitions are smooth in the finite volume system, even at $R$=20 fm, though at such a large volume it is quite a sharp crossover. We also find here a much more enhanced smoothing of the transition with a reduction of volume: while the curve at $R$ = 20 fm indicates a sharp crossover, the one at $R$=5 fm indicates a smooth behavior. Also, the crossover temperature changes much more rapidly here with volume: the transition temperature changes from 47.3 MeV in the infinite volume system to 72.2 MeV in the 5 fm sphere.
@@
Figure 2: The dependence of the $M$ vs $T$ curve with volume. Results at $\mu_{q}$= 100 MeV (left), 250 MeV (middle) and 350 MeV (right) are shown. A discontinuous transition is seen in the infinite volume system at $\mu_{q}$=350 MeV.
Figure 3: The thermodynamic potential, $\Omega/\Lambda_{\rm\scriptscriptstyle UV}^{4}$, in the transition region at $\mu_{q}$=350 MeV for an infinite volume system. A small (irrelevant) additive constant $\Omega=\mp 0.0001\,\Lambda_{\rm\scriptscriptstyle UV}^{4}$ has been added to the potentials at T=44 and 50 MeV, respectively, for better viewing. The figure indicates a first-order transition at $T\approx$ 47.3 MeV.
Before proceeding with the discussion of the investigation of the transition line at higher $\mu_{q}$ at finite volume, let us clarify our terminology. In our analysis, the term ”first-order transition” corresponds to a two-well structure in the free energy, with well-separated local minima at small and large values of $M$ corresponding to the symmetry restored and symmetry broken phases, respectively. Such a structure will indicate phase coexistence, and possible supercooling/superheating in a dynamical system. This is what we have looked for, and have referred to as the signature of a first-order transition.
In the infinite volume case, tunnelling between the two minima is prohibited. However, in a finite volume, the free energy barrier between the minima is finite, and there is a nonzero probability for tunnelling. This will lead to a unique minimum at every $\left(T,\,\mu_{q}\right)$ value for a static system, and strictly speaking, one can only have crossovers in the finite volume system. However, in a fireball which rapidly cools through the transition region, tunnelling will be extremely suppressed. In cases where the classical free energy shows a two-minima structure and a discreet jump at the transition, we will expect interesting phenomena, related to the features of a first-order transition, at the transition point. Therefore, the two-well structure of the mean-field potential, $\Omega_{\scriptscriptstyle MF}$, is of actual physical interest, and we will loosely use terms like first-order line, phase coexistence etc. to refer to such a structure even in a finite volume system.
@@
We illustrate this behavior in Fig. 4, where we show $\Omega_{\scriptscriptstyle MF}$ at different volumes at $T$=20 MeV. We find a two-well structure in $\Omega_{\scriptscriptstyle MF}$ even at $R$=10 fm. Note also the range of $\mu_{q}$ over which this structure is seen (we loosely refer below to this region as the transition region). A shift of the transition to higher $\mu_{q}$ is seen with the reduction in volume; at the same time, a weakening of the double-well structure (and a shrinking of the transition region) is seen. At this temperature, we do not find a discontinuous transition in the 5 fm sphere, but as we go further down to $T$=0 we find a discontinuous transition in all our systems.
Our finding of the transition line in the $(T,\mu_{q})$ plane is summarized in Fig. 5. In the inset, the first-order parts of the line are shown. Here the dotted lines indicate the estimate of the boundary of the two-phase region.
Figure 5: The shift of the chiral crossover line in the $(T,\mu_{q})$ plane with system size. In the inset the large $\mu_{q}$ part is enlarged, highlighting the first-order region.

### v2

@@ Abstract
Using the Nambu Jona-Lasinio model, an effective theory for the chiral transition of quantum chromodynamics (QCD), we have studied the effect of the finite volume of the fireball on the transition line at finite temperature and density using the MIT boundary condition to mimic the condition that the system is deconfined inside. We studied the effect of the finite volume on the transition temperature and on number density and its susceptibilities. The volume effects should be considered when looking for signatures of the phase diagram in experiments.
@@ I Introduction
The bulk property of the strongly interacting matter has been the subject of detailed experimental investigation in recent times [1, 2, 3, 4]. Strongly interacting matter is expected to manifest itself in a deconfined, chiral symmetry-restored phase, called the quark-gluon plasma (QGP), at very high temperatures [5, 6]. Experimentally, this phase has been explored in the study of the medium created in the collision of heavy ions at ultrarelativistic velocities, in particular in the Relativistic Heavy Ion Collider (RHIC) in BNL, USA [3] and the Large Hadron Collider (LHC) in CERN, Geneva [1]. The phase structure at high number density is being explored in the Beam Energy Scan (BES) experiments in RHIC [4] and will be further explored in future experiments in NICA, JINR [7] and in the CBM experiment in FAIR [8]. The knowledge about the phases at high density is complemented by the data from neutron stars and their mergers [9]. On the theory side, the properties of the deconfined state at high temperatures are well-understood from large-scale numerical studies of quantum chromodynamics (QCD), the theory of strongly interacting matter, using lattice discretization [5]. Lattice studies are more difficult at nonzero baryon number chemical potential ($\mu_{\scriptscriptstyle B}$); however, the equation of state and baryon number susceptibilities have been calculated at large temperatures and moderate values of $\mu_{\scriptscriptstyle B}$. At higher values of $\mu_{\scriptscriptstyle B}$, or lower temperatures, one has to rely on QCD-based models [10, 11].
@@
The theoretical studies, in particular the lattice-discretized QCD (LQCD) studies are for static equilibrated infinite volume systems. On the other hand, the medium created in RHIC and LHC experiments is rapidly evolving, and the system size is not too large. In particular, in the BES studies the system dimension can be a few fermi, not too large compared to the typical correlation length in strong interaction physics. The current experimental efforts are to study QGP in smaller systems. All this makes an understanding of the volume effect important.
@@
In this paper, the aim is to get an idea of the effect of the finite volume on the phase diagram for a static system; the effects of the dynamic nature is being left for a future study. We study the phase diagram using the Nambu Jona-Lasinio (NJL) model [10], which is an effective model expected to capture the physics related to the chiral symmetry, in particular, the spontaneous chiral symmetry breaking. Since the finite temperature crossover is associated with the melting of the chiral condensate, the NJL model is expected to capture the basic features of the crossover.
@@
In a study of the finite volume effects it is essential that realistic boundary conditions are put in for the system. The actual boundary of the medium created in the heavy ion collision experiments is of complicated geometry. We take the simplified MIT boundary condition on a spherical geometry [12] (see [13] for a pedagogic discussion). While simple, the MIT boundary condition captures the essential physics that inside the boundary we have a deconfined medium. The spherical MIT bag model has been used to study the effect of the finite size of the QGP droplet on the thermal phase-space distributions of quarks and gluons [14].
@@
The effect of the finite volume on the chiral transition has been studied before [15, 16, 17, 18, 19, 20, 21, 22, 23]. These studies have used the anti-periodic (APBC) and periodic (PBC) boundary conditions, or have put in a lower cutoff in momentum. However, these boundary conditions do not mimic a finite-volume fireball; they are boundary conditions used in theoretical studies to mimic an infinite-volume system within a finite-volume setup, e.g., in numerical lattice studies. In Appendix B we compare the MIT boundary condition results with those obtained with other boundary conditions. The MIT boundary condition has been used before, for studying the chiral crossover at $\mu_{\scriptscriptstyle B}=0$ on a sphere [24], and for a rotating fermionic system in a cylinder [25, 26]. Here we use it to study the phase diagram in the $\left(T,\,\mu_{\scriptscriptstyle B}\right)$ plane and the baryon number susceptibilities.
@@ II Formalism and the boundary condition
The NJL model is an effective model of the spontaneous chiral symmetry breaking of QCD [27, 10]. The degrees of freedom are only quarks; the effects of the gluonic interactions are included in four-fermion interaction terms, chosen such that the Lagrangian is chirally symmetric.
@@
The four-fermion term in $\mathcal{L}_{\scriptscriptstyle\rm NJL}$ makes the theory non-renormalizable. Hence, a suitable regularization method has to be specified, and the results will depend on the regularization parameter. We have used Schwinger’s proper time regularization [27, 28], with the cutoff $\tau_{\rm\scriptscriptstyle UV}=1/\Lambda_{\rm\scriptscriptstyle UV}^{2}$. Therefore, our model has three parameters: $m,G$ and the UV momentum cutoff $\Lambda_{\rm\scriptscriptstyle UV}$. These parameters can be chosen by specifying the pion mass, the pion decay constant and either the chiral condensate or the chiral transition temperature ($T_{\chi}$). Some suitable parameter sets, and how they affect the chiral crossover temperature, have been discussed in Ref. [28]. The phase diagram can be qualitatively different for the different parameter sets. In particular, for some choices of parameters one gets a critical point and a first-order transition at large quark number chemical potential $\mu=\dfrac{\mu_{\scriptscriptstyle B}}{3}$, while for some other choices, the transition is a crossover in the whole $\left(T,\,\mu\right)$ plane. We have chosen the following parameter set:
@@
This parameter set was already explored in Ref. [28]. It has a first-order transition line at large $\mu$, ending at a critical point at $\mu_{c}\approx$ 340 MeV. At $\mu=0$, it shows a chiral transition at $T_{\chi}\approx$ 186 MeV. These values are larger than those expected for QCD: lattice calculations find $T_{\chi}\approx$ 156 MeV[29], and estimates of $\mu_{q,c}$, while less certain, indicate values $<$ 200 MeV [30]. Here our aim is not to tune the absolute value of these parameters but to look at their variation with the system size.
We will be studying the system at finite temperature $T$ and quark number chemical potential $\mu$. $\mu$ is introduced by adding $\mu\psi^{\dagger}\psi$ to Eq. (1). To get the thermodynamic potential, the integral over $p_{0}$ in Eq. (4) has to be turned into a fermionic Matsubara sum. For finite volume, the integral over spatial momenta, $\dfrac{d^{3}p}{(2\pi)^{3}}$, will be replaced by a sum over the momentum modes allowed by the boundary condition. For a study of the effect of finite volume, imposing a proper boundary condition is important. As we discussed in Section I, the exact boundary condition for the medium created in RHIC or LHC is complicated; because it is time-dependent and also changes from event to event. Since we are only interested in getting an idea about how different the finite volume system can be from an infinite volume system, we will take a simple spherical geometry. The important physics that we want to capture is that the deconfined QGP is only contained within this geometry. For this, we impose the MIT boundary condition, which implies that the normal component of the quark current becomes zero at the boundary. More precisely, we impose the boundary condition [12, 13],
@@
Effects of finite volume on the phase diagram have been explored in the literature in the past. Sometimes the antiperiodic or the periodic boundary conditions have been used for this purpose [15, 18, 19, 31, 22]. Note that these boundary conditions do not capture the finite volume fireball in any way: these are boundary conditions usually imposed in theoretical studies (e.g. numerical studies on the lattice) when the aim is to study the infinite volume system. Besides being easy to implement, the advantage of these boundary conditions is that they show very small deviations from the infinite volume system, and are therefore ideal for studies aiming for the thermodynamic limit. Our purpose here is to estimate the size of the deviation from the infinite volume results due to the finite volume fireball. We do not think (anti)periodic boundary conditions are suitable for this purpose. While the boundary condition Eq. (5), Eq. (6) also do not capture the whole complexity of the finite fireball, we believe it captures the most essential aspects of it and gives a more realistic estimation of the size of the finite volume effects. In Appendix B we compare our results with those of an antiperiodic geometry for a similar system size. We indeed find that the volume effects are much smaller in the antiperiodic geometries.
@@ III Chiral transition at finite volume
We explore the chiral transition line for Eq. (3), and its shift with volume, for the MIT boundary condition. At $\mu=0$ the chiral symmetry restoration is a crossover. In Fig. 1, we show the constituent mass, $M$ (Eq. (10)), which is just a constant shift of $-\langle\bar{\psi}\psi\rangle$ and has the same information as $\langle\bar{\psi}\psi\rangle$. We see a rapid but smooth decrease in $M$ around $T\approx$ 190 MeV, indicating a crossover. At low temperatures $\ll T_{\chi}$, decreasing the volume to $R=5$ fm decreases the constituent mass by $<5\%$ from the infinite volume value. At high temperatures $>200$ MeV the chiral symmetry is restored and the chiral condensate becomes small. Therefore $M$ becomes small, approaching $m$. In small-volume systems, the crossover region becomes somewhat smoother, but no qualitative change from the infinite-volume case is seen in systems as small as a 5 fm sphere. This can be further seen in the right panel of Fig. 1, where the derivative $dM/dT$ is shown. To get this plot, we used a cubic spline interpolation of the $M$ vs $T$ data of the left plot. The plot shows a systematic but moderate reduction in the peak height and broadening of the peak, as the volume is reduced. The peak position is an estimate of the crossover temperature $T_{\chi}$, which changes very slowly with the radius of the sphere, $R$ (a $7$ MeV shift from infinite volume to $R=5$ fm). The chiral susceptibility $\chi_{m}=-\frac{\textstyle\partial\langle\bar{\psi}\psi\rangle}{\textstyle\partial m}$ also shows a peak at the chiral symmetry restoration point. The peak position is close to the peak of $dM/dT$, though not exactly identical at the crossover region, and the volume dependence is similar to that seen in Fig. 1. Here we will use the peak of $dM/dT$ as our estimate for $T_{\chi}$. Of course, in the case of a true phase transition (as opposed to a crossover), all these observables will show the same transition temperature.
@@
Figure 1: (Left) Constituent quark mass $M$ as a function of temperature ($T$) at $\mu$ = 0. (Right) The derivative $dM/dT$. The legend shows the peak position of the $dM/dT$ curves, which can be taken as an estimate of $T_{\chi}$.
@@
Now we introduce a nonzero quark chemical potential $\mu$. For small values of $\mu$, the qualitative behavior is very similar to what is seen for $\mu=0$. In Fig. 2, we show $M$ as a function of $T$ at different volumes for three values of $\mu$. The temperature and volume dependence of $M$ at $\mu$=100 MeV is very similar to that at $\mu$=0, as a comparison of the left panel of Fig. 2 with Fig. 1 shows. As $\mu$ increases, the crossover shifts to smaller values of temperature, as expected. Also, we find a stiffening of the transition at larger volumes: the smoothing of the transition at small volumes is enhanced. At $\mu$=350 MeV, shown in the right panel of Fig. 2, we find a discontinuity, indicating a first-order transition, at infinite volume. The first-order nature of the transition can be further ascertained by looking at $\Omega$, Fig. 3. At the transition region, a two-minimum structure is seen, which is the hallmark of first-order transition, and indicates phase coexistence. At this $\mu$ the transitions are smooth in the finite volume system, even at $R$=20 fm, though at such a large volume it is quite a sharp crossover. We also find here a much more enhanced smoothing of the transition with a reduction of volume: while the curve at $R$ = 20 fm indicates a sharp crossover, the one at $R$=5 fm indicates a smooth behavior. Also, the crossover temperature changes much more rapidly here with volume: the transition temperature changes from 47.3 MeV in the infinite volume system to 72.2 MeV in the 5 fm sphere.
@@
Figure 2: The dependence of the $M$ vs $T$ curve with volume. Results at $\mu$=100 MeV (left), 250 MeV (middle) and 350 MeV (right) is shown. A discontinuous transition is seen in the infinite volume system at $\mu$=350 MeV.
Figure 3: The thermodynamic potential, $\Omega/\Lambda_{\rm\scriptscriptstyle UV}^{4}$, in the transition region at $\mu$=350 MeV for an infinite volume system. A small (irrelevant) additive constant $\Omega=\mp 0.0001\,\Lambda_{\rm\scriptscriptstyle UV}^{4}$ has been added to the potentials at T=44 and 50 MeV, respectively, for better viewing. The figure indicates a first-order transition at $T\approx$ 47.3 MeV.
Before proceeding with the discussion of the investigation of the transition line at higher $\mu$ at finite volume, let us clarify our terminology. In our analysis, the term “first-order transition” corresponds to a two-well structure in the free energy, with well-separated local minima at small and large values of $M$ corresponding to the symmetry restored and symmetry broken phases, respectively. Such a structure will indicate phase coexistence, and possible supercooling/superheating in a dynamical system. This is what we have looked for, and have referred to as the signature of a first-order transition.
In the infinite volume case, tunnelling between the two minima is prohibited. However, in a finite volume, the free energy barrier between the minima is finite, and there is a nonzero probability for tunnelling. This will lead to a unique minimum at every $\left(T,\,\mu\right)$ value for a static system, and strictly speaking, one can only have crossovers in the finite volume system. However, the tunnelling probability is exponentially suppressed in volume. In a fireball which rapidly cools through the transition region, tunnelling will be insignificant unless the volume is very small. In cases where the classical free energy shows a two-minima structure and a discreet jump at the transition, we will expect interesting phenomena, related to the features of a first-order transition, at the transition point (see also Ref.[15]). Therefore, the two-well structure of the mean-field potential, $\Omega$, is of actual physical interest, and we will loosely use terms like first-order line, phase coexistence etc. to refer to such a structure even in a finite volume system. An analysis of the dynamics at discontinuity in a finite volume system has been done in Ref. [32].
@@
We illustrate this behavior in Fig. 4, where we show $\Omega$ at different volumes at $T$=20 MeV. We find a two-well structure in $\Omega$ even at $R$=10 fm. Note also the range of $\mu$ over which this structure is seen (we loosely refer below to this region as the coexistence region). A shift of the transition to higher $\mu$ is seen with the reduction in volume; at the same time, a weakening of the double-well structure (and a shrinking of the coexistence region) is seen. At this temperature, we do not find a discontinuous transition in the 5 fm sphere, but as we go further down to $T$=0 we find a discontinuous transition in all our systems.
Our finding of the transition line in the $\left(T,\,\mu\right)$ plane is summarized in Fig. 5. The dashed line corresponds to the smooth crossover region, whereas the solid part of the line at small $T$ indicates the region where the mean field effective potential behaves like a first order transition. We find a shift of the crossover line towards larger values of $\mu$ as the system size is reduced. The shift is mild for systems with radii as small as 10 fm, but a more significant shift is seen for the 5 fm sphere.
A more significant volume dependence is found when we investigate the first order transition region at low temperatures. This part is shown more clearly in the inset. As the figure shows, the coexistence region in the mean field potential is pushed down to lower values of $T$ as the system radius is decreased. The dotted lines around the first order line in the inset indicate the estimate of the boundary of the coexistence region. This region shrinks quite strongly as the system size is reduced.
Figure 5: The shift of the chiral crossover line in the $(T,\mu)$ plane with system size. In the inset the large $\mu$ part is enlarged, highlighting the first-order region.